OpenStudy (anonymous):

which of the following is a correct factorization of this trinomial? -3x^2+10x-8 A. (-3x+4)(x-2) B. -(3x+4)(x+2) C. -(x+4)(x-3) D. -3(x+4)(x+2)

OpenStudy (anonymous):

-(x-2) (3 x-4) A.

OpenStudy (er.mohd.amir):

-3x*x+10x-8 =-3x*x+6x+4x-8 =-3x(x-2)+4(x-2) =(x-2)(-3x+4)